# Semi-Supervised Clustering of Sparse Graphs: Crossing the Information-Theoretic Threshold

### 08 Jan 2024 v3

@@ 4 Semi-Supervised SDP
We have seen that the census method solves the semi-supervised community detection problem. But the algorithm is desirable in practice only when the amount of revealed labels is sufficient to support a reasonable performance. In other words, it has no unsupervised ‘fallback’ built in. Meanwhile, SDPs enjoy nice properties like optimality and robustness as mentioned earlier. It is also well known that approximate information about the extremal cuts of graphs can be obtained by computing the optimizer for SDP of their adjacency matrix, see for example (Goemans and Williamson, 1995). From both a practical and a mathematical point of view, we are interested in developing an SDP based semi-supervised clustering approach, and through which we shall be able to see the models, algorithms and phase transitions with a fresh perspective.
In this section, we will focus on the hypothesis testing formulation of the community detection problem. We have discussed the equivalency between it and the non-vanishing overlap formulation under the unsupervised setting. In the semi-supervised scenario it is still an interesting question to ask whether there exists a test that can distinguish SBMs from ERMs. Here we understand ERM as the special case of SBM with $a=b$. It also has ground truth of labels, which is uniformly random under the balance constraint. Given that they are originally contiguous when $\text{SNR}\leq 1$, we want to show that revealed labels together with random graphs can separate them.
@@ 4.1 SDP for Community Detection
Under the Planted Bisection Model specified in Definition 1, MAP estimator is equivalent to the Maximum Likelihood estimator, which is given by min-bisection, i.e., a balanced partition with the least number of crossing edges. Formally, it can be written the following optimization problem,
$\max_{\begin{subarray}{c}x\in\{1,-1\}^{n}\\ x^{\top}\mathbf{1}=0\end{subarray}}x^{\top}Ax$ (65)
@@
$\hat{X}_{\text{MAP}}(G)=\operatorname*{arg\,max}_{\begin{subarray}{c}X\succeq 0\\ X_{ii}=1,~{}\forall i\in[n]\\ \text{rank}(X)=1\\ X\mathbf{1}=0\end{subarray}}\langle A,X\rangle$ (66)
Although min-bisection of $G$ is optimal (in the MAP sense) for exact recovery, finding it is NP-hard. Various relaxations have been proposed for the MAP estimator. Since the rank constraint makes the optimization difficult, we can remove it to make the problem convex. One can also get rid of the balance constraint by centralizing the adjacency matrix, $\tilde{A}\coloneqq A-\frac{d}{n}\mathbf{1}\mathbf{1}^{\top}$ with $d=(a+b)/2$ the average degree. This can also be justified using Lagrangian multipliers. And the resulting semidefinite relaxation is given by
$\hat{X}_{\text{SDP}}(G)=\operatorname*{arg\,max}_{\begin{subarray}{c}X\succeq 0\\ X_{ii}=1,~{}\forall i\in[n]\end{subarray}}\langle\tilde{A},X\rangle$ (67)
@@
$\text{SDP}(M_{n\times n})=\max\{\langle M,X\rangle\mathrel{\mathop{:}}X\in\text{elliptope}_{n}\}$ (68)
@@
Proof  For any feasible $X\succeq 0$ and $X_{ii}=1$, we have $\operatorname{Tr}(X)=n$.
$\langle X,M=U\Lambda U^{\top}\rangle=\operatorname{Tr}(U^{\top}XU\Lambda)=\langle Y\coloneqq U^{\top}XU,\Lambda\rangle=\sum Y_{ii}\lambda_{i}\leq n\lambda_{1}$ (70)
@@
This proposition relates SDPs to spectra of the underlying matrices, which suffer from those high degree nodes as we mentioned in the introduction. In contrast, SDPs behave similarly on SBMs and random regular graphs. The optimal values of the SDPs for both are approximately $2n\sqrt{d}$, see (Montanari and Sen, 2016). Random regular graphs obey the uniform distribution over graphs with $n$ vertices and uniform degree $d$, which provide a simple example to illustrate the regularity property of SDPs. We cite an intermediate result from the original proof as Lemma 25.
An important way to understand SDPs is considering the Burer-Monteiro factorization of $X$, which characterizes the constraints. We have
@@
where $\Sigma=(\sigma_{1},\sigma_{2},\dots,\sigma_{n})^{\top}$ and $\|\sigma_{i}\|_{2}=1,~{}\forall i\in[n]$. Therefore, the $i$-th node of the graph is associated with the vector $\sigma_{i}$ that lies on the unit sphere. $X_{ij}=\langle\sigma_{i},\sigma_{j}\rangle$ can be interpreted as the affinity metric between nodes $i$ and $j$. SDP maximizes the likelihood score of this affinity matrix with respect to the given centralized adjacency matrix. The optimizer $X^{*}$ is a better representation of the structure information than the vanilla adjacency matrix. Then we can identify the labels by simply running the k-means method on it or compute the eigenvector corresponding to the largest eigenvalue.
@@ 4.2 Constrained SDP and Semi-Supervised Testing
In this section, we introduce our SDP modification and prove that it solves the semi-supervised community detection problem with the hypothesis testing formulation. Let $x$ denote labels of G(n, $\frac{a}{n}$, $\frac{b}{n}$). And $m$ of them are revealed uniformly at random in a balanced manner. Conditioning on the ground truth of clusters, indices of revealed nodes $\mathcal{R}$ and edges are independent. So without loss of generality, we denote revealed labels $\tilde{x}$ as follows.
$x_{i}=\begin{cases}1,\quad&i=1,2,\dots,\frac{n}{2}\\ -1,\quad&i=\frac{n}{2},\frac{n}{2}+1,\dots,n\end{cases}\quad\tilde{x}_{i}=\begin{cases}1,\quad&i=1,2,\dots,\frac{m}{2}\\ -1,\quad&i=\frac{n}{2},\frac{n}{2}+1,\dots,\frac{n+m}{2}\\ 0,\quad&\text{otherwise}\end{cases}$ (72)
We have shown that the entry value of the optimizer $X$ can be interpreted as an affinity metric among nodes. Moreover, we have $X_{ij}\in[-1,1],~{}\forall~{}i,j$. It is natural to force the optimizer to have large entry values for those vertex pairs in which we have high confidence to be in the same community and vice versa. Therefore, we propose the CSDP approach to integrate the information provided by semi-supervised approach. If node i and node j are revealed to have the same label, we add the constraint $X_{ij}=1$ to the optimization model. If they are revealed to have the opposite labels, we add $X_{ij}=-1$. Formally, the CSDP is defined as
@@
where $\mathcal{R}$ denotes the collection of revealed nodes. After reordering the indices, we can assume it as $\{1,2,\dots,\frac{m}{2}\}\bigcup\{\frac{n}{2},\frac{n}{2}+1,\dots,\frac{n+m}{2}\}$. It is worth noting that the optimization remains a positive semidefinite programming, which can be solved efficiently, for example by interior point methods (Alizadeh, 1995).
Then let $\mathcal{S}^{n-1}\coloneqq\{v\in\mathbb{R}^{n}\mathrel{\mathop{:}}\|v\|_{2}=1\}$ be the unit $(n-1)$-sphere and $\sigma=(\sigma_{1},\sigma_{2},\dots,\sigma_{n})\in(\mathcal{S}^{n-1})^{n}$. Consider the CSDP in the form of Burer-Monteiro factorization. We have the identity
$\text{SDP}(M)=\max\left\{\sum_{i,j=1}^{n}M_{ij}\langle\sigma_{i},\sigma_{j}\rangle\mathrel{\mathop{:}}\sigma_{i}\in\mathcal{S}^{n-1}~{}\forall i\in[n]\right\}$ (74)
@@
$M^{\text{agg}}_{00}$ $=\sum_{i,j\in\mathcal{R}}M_{ij}x_{i}x_{j}$ (77)
$M^{\text{agg}}_{0j}$ $=\sum_{i\in\mathcal{R}}x_{i}M_{i,j+\frac{m}{2}},~{}\forall j\in[\frac{n}{2}-\frac{m}{2}]$ (78)
$M^{\text{agg}}_{0j}$ $=\sum_{i\in\mathcal{R}}x_{i}M_{i,j+m},~{}\forall j\in\in[n-m]\setminus[\frac{n}{2}-\frac{m}{2}]$ (79)
$M^{\text{agg}}_{ij}$ $=M_{i+\frac{m}{2},j+\frac{m}{2}},~{}\forall i,j\in[\frac{n}{2}-\frac{m}{2}]$ (80)
$M^{\text{agg}}_{ij}$ $=M_{i+m,j+m},~{}\forall i,j\in[n-m]\setminus[\frac{n}{2}-\frac{m}{2}]$ (81)
Essentially, we aggregate the rows and columns related to revealed vertices according to their communities into the 00-th row and column and reindex the matrix. It introduces spikiness to the underlying matrix.
$M^{\text{agg}}=\left(\begin{array}[]{@{}c|c@{}}\sum_{i,j\in\mathcal{R}}M_{ij}x_{i}x_{j}&\begin{matrix}M^{\text{agg}}_{01}&M^{\text{agg}}_{02}&\cdots&M^{\text{agg}}_{0,n-m}\end{matrix}\\ \hline\cr\begin{matrix}M^{\text{agg}}_{01}\\ M^{\text{agg}}_{02}\\ \vdots\\ M^{\text{agg}}_{0,n-m}\end{matrix}&M_{\mathcal{R^{\complement}}}\end{array}\right)$ (82)
Although (Montanari and Sen, 2016) takes a rather different approach to study SDPs, they also notice that the critical change comes with such built-in structures, where the authors state ”we expect the phase transition in $\text{SDP}(\lambda vv^{\top}+W)/n$ to depend—in general—on the vector $v$, and in particular on how ‘spiky’ this is”.
@@
$\hat{X}^{*}_{ij}=\begin{cases}1&i=j=0\\ 0&~{}i\in[n-m],~{}j=0\\ 0&~{}j\in[n-m],~{}i=0\\ X^{*}_{ij}&~{}\text{otherwise}\\ \end{cases}$ (88)
@@
$\text{CSDP}(M)=\text{SDP}(M^{\text{agg}})\geq\langle\hat{X^{*}},M^{\text{agg}}\rangle=\text{SDP}(M_{\mathcal{R}^{\complement}})+M^{\text{agg}}_{00}$ (89)
So far, all the results are deterministic, $M$ can be arbitrary symmetric matrix and $\mathcal{R}$ can be any balanced index set. Next, we will consider $M=\tilde{A}\coloneqq A-\frac{d}{n}\mathbf{1}\mathbf{1}^{\top}$ to study CSDPs on probabilistic models.
@@
The CSDP always lies in between the SDPs of the original adjacency matrix and the submatrix of unrevealed vertices. Moreover, if $\tilde{A}\sim\operatorname{\mathcal{G}}(n,\frac{a}{n},\frac{b}{n})$, we have $\tilde{A}_{\mathcal{R}^{\complement}}\sim\operatorname{\mathcal{G}}(n-m,\frac{a(1-\rho)}{n-m},\frac{b(1-\rho)}{n-m})$. It is worth mentioning that although $\tilde{A}_{\mathcal{R}^{\complement}}$ is just a submatrix of the original centered adjacency matrix, its probabilistic distribution as a random matrix is not simply changed from $n$ nodes to $n-m$ nodes. The edge probability parameters are also changed by a factor of $(1-\rho)$. It leads to some technical challenges, which we are going to handle later. But intuitively, from the asymptotic behavior of SDP, we can derive a rough understanding of CSDP as $n\to\infty$. Recall that the phase transition theory tells us that when SNR $\leq$ 1, the optimal value of SDP for SBM will not be large enough to distinguish from the optimal value of SDP for ERM. Therefore, the order of above quantities from inequality (90) suggests that semi-supervised SDP can not help to increase the statistics associated with SBM. The best one can hope for is that it will make the statistics associated with ERM smaller by a factor depending on $\rho$. This turns out to be enough for community detection.
Recall the community detection problem can be formalized as a binary hypothesis testing problem, whereby we want to determine, with high probability of success, whether the random graph under consideration has a community structure or not. As discussed in Section 2, we introduce semi-supervised learning to the problem by revealing a part of the labels involved in the random graph generating process. Namely, if the labels associated with a graph $G$ over $n$ vertices are denoted as $x$, we choose $m$ of them uniformly at random denote the index set by $\mathcal{R}$, such that $\sum_{i\in\mathcal{R}}x_{i}=0$.
@@
A statistical test $T$ is a function defined on the graphs and revealed labels with range $\{0,1\}$. It succeeds with high probability if
$\operatorname{P}_{0}(T(G,x_{\mathcal{R}})=1)+\operatorname{P}_{1}(T(G,x_{\mathcal{R}})=0)\to 0\quad(n\to\infty)$ (91)
Notice that this is indeed a generalization of the unsupervised community detection. Simply looking into the labels, two models are indistinguishable. What characters their difference is the probabilistic law of how edges are generated, i.e., whether there is a cluster structure. The revealed labels serve as an enhancement of the graph observed. The phase transition theory says that under the unsupervised setting (the special case when $\rho=0$), no test can succeed with high probability when SNR $\leq 1$, or equivalently, $a-b\leq\sqrt{2(a+b)}$. While if $\text{SNR}>1$, several polynomially computable tests are developed. SDP based test is nearly optimal, in the sense that it requires

### 28 Feb 2024 v4

@@ 4 Semi-Supervised SDP
We have seen that the census method solves the semi-supervised community detection problem. However, the algorithm is desirable in practice only when the amount of revealed labels is sufficient to support a reasonable performance. In other words, it has no unsupervised ‘fallback’ built-in. Meanwhile, SDPs enjoy nice properties like optimality and robustness as mentioned earlier. It is also well known that approximate information about the extremal cuts of graphs can be obtained by computing the optimizer for the SDP of their adjacency matrix, for example (Goemans and Williamson, 1995). From both a practical and a mathematical point of view, we are interested in developing an SDP-based semi-supervised clustering approach, through which we shall be able to see the models, algorithms, and phase transitions with a fresh perspective.
In this section, we will focus on the hypothesis testing formulation of the community detection problem. We have discussed the equivalency between it and the non-vanishing overlap formulation under the unsupervised setting. In the semi-supervised scenario, it is still an interesting question to ask whether there exists a test that can distinguish SBMs from ERMs. Here we understand ERM as the special case of SBM with $a=b$. It also has ground truth of labels, which is uniformly random under the balance constraint. Given that they are originally contiguous when $\text{SNR}\leq 1$, we want to show that revealed labels together with random graphs can separate them.
@@ 4.1 SDP for Community Detection
Under the Planted Bisection Model specified in Definition 1, the MAP estimator is equivalent to the Maximum Likelihood estimator, which is given by min-bisection, i.e., a balanced partition with the least number of crossing edges. Formally, it can be written the following optimization problem,
$\max_{\begin{subarray}{c}x\in\{1,-1\}^{n}\\ x^{\top}\mathbf{1}=0\end{subarray}}x^{\top}Ax.$ (65)
@@
$\hat{X}_{\text{MAP}}(G)=\operatorname*{arg\,max}_{\begin{subarray}{c}X\succeq 0\\ X_{ii}=1,~{}\forall i\in[n]\\ \text{rank}(X)=1\\ X\mathbf{1}=0\end{subarray}}\langle A,X\rangle.$ (66)
Although min-bisection of $G$ is optimal (in the MAP sense) for exact recovery, finding it is NP-hard. Various relaxations have been proposed for the MAP estimator. Since the rank constraint makes the optimization difficult, we can remove it to make the problem convex. One can also get rid of the balance constraint by centralizing the adjacency matrix, $\tilde{A}\coloneqq A-\frac{d}{n}\mathbf{1}\mathbf{1}^{\top}$ with $d=(a+b)/2$ the average degree. This can also be justified using Lagrangian multipliers. The resulting semidefinite relaxation is given by
$\hat{X}_{\text{SDP}}(G)=\operatorname*{arg\,max}_{\begin{subarray}{c}X\succeq 0\\ X_{ii}=1,~{}\forall i\in[n]\end{subarray}}\langle\tilde{A},X\rangle.$ (67)
@@
$\text{SDP}(M_{n\times n})=\max\{\langle M,X\rangle\mathrel{\mathop{:}}X\in\text{elliptope}_{n}\}.$ (68)
@@
Proof  For any feasible $X\succeq 0$ and $X_{ii}=1$, we have $\operatorname{Tr}(X)=n$, moreover,
$\langle X,M=U\Lambda U^{\top}\rangle=\operatorname{Tr}(U^{\top}XU\Lambda)=\langle Y\coloneqq U^{\top}XU,\Lambda\rangle=\sum Y_{ii}\lambda_{i}\leq n\lambda_{1}.$ (70)
@@
This proposition relates SDPs to spectra of the underlying matrices, which suffer from those high-degree nodes as we mentioned in the introduction. In contrast, SDPs behave similarly on SBMs and random regular graphs. The optimal values of the SDPs for both are approximately $2n\sqrt{d}$, see (Montanari and Sen, 2016). Random regular graphs obey the uniform distribution over graphs with $n$ vertices and uniform degree $d$, which provide a simple example to illustrate the regularity property of SDPs. We cite an intermediate result from the original proof as Lemma 25.
An important way to understand SDPs is by considering the Burer-Monteiro factorization of $X$, which characterizes the constraints. We have
@@
where $\Sigma=(\sigma_{1},\sigma_{2},\dots,\sigma_{n})^{\top}$ and $\|\sigma_{i}\|_{2}=1,~{}\forall i\in[n]$. Therefore, the $i$-th node of the graph is associated with the vector $\sigma_{i}$ that lies on the unit sphere. $X_{ij}=\langle\sigma_{i},\sigma_{j}\rangle$ can be interpreted as the affinity metric between nodes $i$ and $j$. SDP maximizes the likelihood score of this affinity matrix concerning the given centralized adjacency matrix. The optimizer $X^{*}$ is a better representation of the structure information than the vanilla adjacency matrix. Then we can identify the labels by simply running the k-means method on it or compute the eigenvector corresponding to the largest eigenvalue.
@@ 4.2 Constrained SDP and Semi-Supervised Testing
In this section, we will introduce our SDP modification and prove that it solves the semi-supervised community detection problem with the hypothesis testing formulation. Let $x$ denote labels of G(n, $\frac{a}{n}$, $\frac{b}{n}$). And $m$ of them are revealed uniformly at random in a balanced manner. Conditioned on the ground truth of clusters, indices of revealed nodes $\mathcal{R}$ and edges are independent. So, without loss of generality, we denote revealed labels $\tilde{x}$ as follows:
$x_{i}=\begin{cases}1\quad&i=1,2,\dots,\frac{n}{2},\\ -1\quad&i=\frac{n}{2},\frac{n}{2}+1,\dots,n;\end{cases}\quad\tilde{x}_{i}=\begin{cases}1\quad&i=1,2,\dots,\frac{m}{2},\\ -1\quad&i=\frac{n}{2},\frac{n}{2}+1,\dots,\frac{n+m}{2},\\ 0\quad&\text{otherwise}.\end{cases}$ (72)
We have shown that the entry value of the optimizer $X$ can be interpreted as an affinity metric among nodes. Moreover, we have $X_{ij}\in[-1,1],~{}\forall~{}i,j$. It is natural to force the optimizer to have large entry values for those vertex pairs in which we have high confidence to be in the same community and vice versa. Therefore, we propose the CSDP approach to integrate the information provided by the semi-supervised approach. If node i and node j are revealed to have the same label, we add the constraint $X_{ij}=1$ to the optimization model. If they are revealed to have the opposite labels, we add $X_{ij}=-1$. Formally, the CSDP is defined as
@@
where $\mathcal{R}$ denotes the collection of revealed nodes. After reordering the indices, we can assume it as $\{1,2,\dots,\frac{m}{2}\}\bigcup\{\frac{n}{2},\frac{n}{2}+1,\dots,\frac{n+m}{2}\}$. It is worth noting that the optimization remains a positive semidefinite programming problem, which can be solved efficiently, for example by interior point methods (Alizadeh, 1995).
Then let $\mathcal{S}^{n-1}\coloneqq\{v\in\mathbb{R}^{n}\mathrel{\mathop{:}}\|v\|_{2}=1\}$ be the unit $(n-1)$-sphere and $\sigma=(\sigma_{1},\sigma_{2},\dots,\sigma_{n})\in(\mathcal{S}^{n-1})^{n}$. Consider the CSDP in the form of Burer-Monteiro factorization. We have the following identities:
$\text{SDP}(M)=\max\left\{\sum_{i,j=1}^{n}M_{ij}\langle\sigma_{i},\sigma_{j}\rangle\mathrel{\mathop{:}}\sigma_{i}\in\mathcal{S}^{n-1}~{}\forall i\in[n]\right\},$ (74)
@@
$M^{\text{agg}}_{00}$ $=\sum_{i,j\in\mathcal{R}}M_{ij}x_{i}x_{j},$ (77)
$M^{\text{agg}}_{0j}$ $=\sum_{i\in\mathcal{R}}x_{i}M_{i,j+\frac{m}{2}}\qquad\forall j\in[\frac{n}{2}-\frac{m}{2}],$ (78)
$M^{\text{agg}}_{0j}$ $=\sum_{i\in\mathcal{R}}x_{i}M_{i,j+m}\qquad\forall j\in[n-m]\setminus[\frac{n}{2}-\frac{m}{2}],$ (79)
$M^{\text{agg}}_{ij}$ $=M_{i+\frac{m}{2},j+\frac{m}{2}}\qquad\forall i,j\in[\frac{n}{2}-\frac{m}{2}],$ (80)
$M^{\text{agg}}_{ij}$ $=M_{i+m,j+m}\qquad\forall i,j\in[n-m]\setminus[\frac{n}{2}-\frac{m}{2}].$ (81)
Essentially, we aggregate the rows and columns related to revealed vertices according to their communities into the 00-th row and column and reindex the matrix. It introduces spikiness to the underlying matrix,
$M^{\text{agg}}=\left(\begin{array}[]{@{}c|c@{}}\sum_{i,j\in\mathcal{R}}M_{ij}x_{i}x_{j}&\begin{matrix}M^{\text{agg}}_{01}&M^{\text{agg}}_{02}&\cdots&M^{\text{agg}}_{0,n-m}\end{matrix}\\ \hline\cr\begin{matrix}M^{\text{agg}}_{01}\\ M^{\text{agg}}_{02}\\ \vdots\\ M^{\text{agg}}_{0,n-m}\end{matrix}&M_{\mathcal{R^{\complement}}}\end{array}\right).$ (82)
Although (Montanari and Sen, 2016) employed a rather different technique to study SDPs, they also noticed that the critical change comes with such built-in structures, where the authors state ”We expect the phase transition in $\text{SDP}(\lambda vv^{\top}+W)/n$ to depend—in general—on the vector $v$, and in particular on how ‘spiky’ this is”.
@@
$\hat{X}^{*}_{ij}=\begin{cases}1&i=j=0,\\ 0&~{}i\in[n-m],~{}j=0,\\ 0&~{}j\in[n-m],~{}i=0,\\ X^{*}_{ij}&~{}\text{otherwise}.\\ \end{cases}$ (88)
@@
$\text{CSDP}(M)=\text{SDP}(M^{\text{agg}})\geq\langle\hat{X^{*}},M^{\text{agg}}\rangle=\text{SDP}(M_{\mathcal{R}^{\complement}})+M^{\text{agg}}_{00}.$ (89)
So far, all the results are deterministic, $M$ can be an arbitrary symmetric matrix, and $\mathcal{R}$ can be any balanced index set. Next, we will consider $M=\tilde{A}\coloneqq A-\frac{d}{n}\mathbf{1}\mathbf{1}^{\top}$ to study CSDPs on probabilistic models.
@@
The CSDP always lies in between the SDPs of the original adjacency matrix and the submatrix of unrevealed vertices. Moreover, if $\tilde{A}\sim\operatorname{\mathcal{G}}(n,\frac{a}{n},\frac{b}{n})$, we have $\tilde{A}_{\mathcal{R}^{\complement}}\sim\operatorname{\mathcal{G}}(n-m,\frac{a(1-\rho)}{n-m},\frac{b(1-\rho)}{n-m})$. It is worth mentioning that although $\tilde{A}_{\mathcal{R}^{\complement}}$ is just a submatrix of the original centered adjacency matrix, its probabilistic distribution as a random matrix is not simply changed from $n$ nodes to $n-m$ nodes. The edge probability parameters are also changed by a factor of $(1-\rho)$. It leads to some technical challenges, which we are going to handle later. But intuitively, from the asymptotic behavior of SDP, we can derive a rough understanding of CSDP as $n\to\infty$. Recall that the phase transition theory tells us that when SNR $\leq$ 1, the optimal value of SDP for SBM will not be large enough to distinguish from the optimal value of SDP for ERM. Therefore, the order of the above quantities from inequality (90) suggests that semi-supervised SDP can not help to increase the statistics associated with SBM. The best one can hope for is that it will make the statistics associated with ERM smaller by a factor depending on $\rho$. This turns out to be enough for community detection.
Recall the community detection problem can be formalized as a binary hypothesis testing problem, whereby we want to determine, with a high probability of success, whether the random graph under consideration has a community structure or not. As discussed in Section 2, we introduce semi-supervised learning to the problem by revealing a part of the labels involved in the random graph’s generating process. Namely, if the labels associated with a graph $G$ over $n$ vertices are denoted as $x$, we choose $m$ of them uniformly at random and denote the index set by $\mathcal{R}$, such that $\sum_{i\in\mathcal{R}}x_{i}=0$.
@@
A statistical test $T$ is a function defined on the graphs and revealed labels with range $\{0,1\}$. It succeeds with a high probability if
$\operatorname{P}_{0}(T(G,x_{\mathcal{R}})=1)+\operatorname{P}_{1}(T(G,x_{\mathcal{R}})=0)\to 0\quad(n\to\infty).$ (91)
Note that this is indeed a generalization of the unsupervised community detection. Simply looking at the labels, the two models are indistinguishable. What characterizes their difference is the probabilistic law of how edges are generated, i.e., whether there is a cluster structure. The revealed labels serve as an enhancement of the graph observed. The phase transition theory says that under the unsupervised setting (the special case when $\rho=0$), no test can succeed with high probability when SNR $\leq 1$, or equivalently, $a-b\leq\sqrt{2(a+b)}$. While if $\text{SNR}>1$, several polynomially computable tests are developed. The SDP-based test is nearly optimal, in the sense that it requires